OpenStudy (anonymous):

PLEASE HELP --> I WILL GIVE A MEDAL IF YOU EXPLAIN STEP BY STEP FOR EACH PART -------------------------------------------------------------------- Task 3 Part 1: Create a scenario for an arithmetic sequence. For example: Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? (Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation.) __________________________________________________________ Part 2: Create a scenario for a geometric sequence. For example: Anthony goes to the gym for ______ minutes on Monday. Every day he _________his gym time by ____________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? __________________________________________________________ Part 3: Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (benlindquist):

k

OpenStudy (anonymous):

@Nicki15

OpenStudy (anonymous):

@Mehek14

OpenStudy (owlcoffee):

An arithmetic series is a progression of sums of number which behave in a modelled pattern, for example: \[S_n=1+2+3+4+5...+n\] The sum of this is infinite, since there is no visible convergence of the resulting terms. The geometric series is the same an arithmetic series, but each term is multiplied by a ratio: \[S_n=1+\frac{ 1 }{ 2 }+\frac{ 1 }{ 4 }+\frac{ 1 }{ 8 }+...\frac{ 1 }{ 2^n }\] which is this case is 1/2

OpenStudy (anonymous):

*nods head*